# How Efficient Are Diffusion Language Models? A Critical Examination of Efficiency Evaluation Practices

### v1

@@ 5.2.2 Analysis of AR Models
In the decoding phase of AR models, the model generates only one token at each step, incurring per-token FLOPs of $\mathcal{O}(Bd^{2})+\mathcal{O}(BLd)$ and MOPs of $\mathcal{O}(d^{2})+\mathcal{O}(BLd)$ due to Key-Value (KV) cache reads.
According to the Roofline Model, when the batch size is small, the arithmetic intensity of the AR decoding process is relatively low, making it a memory-bound workload. Consequently, its achievable throughput is mainly limited by memory bandwidth and can be asymptotically expressed as:
$\text{Throughput}\approx\begin{cases}\dfrac{\mathcal{O}(B)}{\mathcal{O}(d^{2})+\mathcal{O}(Ld)},&L_{p}\ll d,\\[10.0pt] \dfrac{\mathcal{O}(1)}{\mathcal{O}(d^{2})+\mathcal{O}(Ld)},&L_{p}\gg d.\end{cases}$ (5)
@@
Effect of Batch Size. In our evaluation setting (with generation length fixed at 256 tokens and relatively short prompts), the condition $L_{p}\ll d$ holds. In this regime, the denominator remains constant while the numerator scales linearly with batch size $B$ (because of the arithmetic intensity is $\mathcal{O}(B)$ ), leading to nearly linear throughput growth until the system reaches the compute roof. Once the arithmetic intensity reaches the compute roof, throughput remains constant as the workload becomes compute-bound.
@@ 5.2.3 Analysis of Diffusion Models
Therefore, the arithmetic intensity satisfies
$\text{ArInt}\approx\begin{cases}\mathcal{O}(BL),&L_{p}\ll d,\\[6.0pt] \mathcal{O}(L),&L_{p}\gg d.\end{cases}$ (7)
Effect of Generation Length. For very short $L$ , the workload is still memory-bound. The throughput can be written as
$\text{Throughput}\approx\frac{\mathcal{O}(L)}{\mathcal{O}(Ld^{2})+\mathcal{O}(L^{2}d)}=\frac{\mathcal{O}(1)}{\mathcal{O}(d^{2})+\mathcal{O}(Ld)}.$ (8)
Since $L\ll d$, the $\mathcal{O}(d^{2})$ term dominates, so the throughput is roughly stable with respect to $L$.
In most practical settings, the DLM is compute-bound. In this case, the throughput is
$\text{Throughput}\approx\frac{P_{max}}{\mathcal{O}(Ld^{2})+\mathcal{O}(L^{2}d)}.$ (9)
Here the numerator stays roughly constant, so as $L$ increases, the overall throughput decreases.
Effect of Batch Size. In our experiments, the diffusion model is compute-bound (with $\text{ArInt}\approx\mathcal{O}(BL)$ ). From the formula above, throughput stays roughly flat as batch size $B$ increases, until it hits the GPU memory limit. This matches the observations in RQ1.
@@ 5.2.4 Analysis of Block Diffusion Models
In the decoding phase of block diffusion models, the sequence is partitioned into blocks of size G. Tokens within each block are generated in parallel via a diffusion process, incurring FLOPs of $\mathcal{O}(BGd^{2})+\mathcal{O}(BGLd)$. Despite the inter-block parallelism, the block-level progression follows an autoregressive scheme, benefiting from KV cache reuse. The resulting arithmetic intensity is inherently higher than that of standard AR models because the expensive KV cache reads are amortized across the parallel block generation, its throughput can be formulated as:
$\text{Throughput}\approx\begin{cases}\dfrac{\mathcal{O}(BG)}{\mathcal{O}(Gd^{2})+\mathcal{O}(GLd)},&L_{p}\ll d,\\[10.0pt] \dfrac{\mathcal{O}(G)}{\mathcal{O}(Gd^{2})+\mathcal{O}(GLd)},&L_{p}\gg d.\end{cases}$ (10)
Effect of Generation Length. Similar to AR, the block diffusion model is memory-bound when the batch size is 1. After simplification, its throughput is the same as that of the AR model, so the throughput trend is identical——roughly constant for short $L$ (dominated by $\mathcal{O}(d^{2})$ ) and decreasing as $L$ grows (dominated by $\mathcal{O}(Ld)$).
Effect of Batch Size. In our evaluation setting (with generation length fixed at 256 tokens and relatively short prompts), the condition $L_{p}\ll d$ holds. For block diffusion models, the arithmetic intensity is $\mathcal{O}(BG)$, which is higher than AR under the same conditions. As a result, block diffusion reaches the compute-bound regime at a smaller batch size than AR. As batch size increases, the throughput of block diffusion models behaves much like AR: throughput grows almost linearly with $B$ until it hits the compute roof; beyond that point, throughput stays steady as the workload becomes compute-bound.
@@ 5.3.1 Effect of sequence length
For block diffusion models, the “+parallel” strategy is highly effective when evaluated with a batch size of 1, consistently achieving the fastest decoding performance across all sequence lengths and even outperforming the AR baseline. We observe that this acceleration effect further amplifies as sequence length increases, reaching up to 3.1× speedup over the original configuration (i.e., 142.33 vs. 46.20 tokens/s). This improvement stems from the model’s ability to generate multiple high-confidence tokens simultaneously within each forward pass, effectively reducing the number of decoding steps and increasing overall throughput. Consequently, the effectiveness of different parallel strategies can be evaluated by their degree of parallel generation, which can be quantified by the number of tokens per forward step (TPF) [18]. This metric directly reflects how efficiently a model converts computational intensity into output throughput, providing a unified lens for analyzing diffusion-based acceleration methods.
@@ 6 Future Directions
Inference Infrastructure for DLMs. Despite the impressive progress of DLMs, major machine learning ecosystems provide only limited optimization and deployment support for DLMs, posing practical challenges for researchers and developers. In particular, DLMs lack mature, open-source inference infrastructure akin to vLLM [19], making efficient serving of DLMs difficult.
@@
Applications of DLMs. Although DLMs are generally less efficient than AR models, their parallel architecture and iterative refinement can offer advantages in compute-rich or latency-sensitive scenarios. We highlight two representative application directions: (1) on-device inference. Combining with parallel decoding techniques, DLMs can achieve higher arithmetic intensity and throughput efficiency, making them well-suited for latency-sensitive and low-concurrency environments such as in-vehicle systems, mobile devices, and edge computing platforms. (2) structured and non-sequential generation. DLMs exhibit notable advantages in structured or non-sequential generation tasks such as code synthesis, table generation, and logical planning [20, 21]. This suggests that the DLMs hold strong potential for scenarios requiring global semantic planning or multi-constraint generation.
@@ 7 Conclusion
In this report, we present a preliminary investigation into the efficiency of DLMs through both theoretical and empirical analyses. Our analysis reveals that existing evaluations remain limited in scope and consistency, lacking a comprehensive understanding of DLM efficiency across diverse conditions. In response, we introduce a theoretical analysis based on the roofline model, revealing that while DLMs exhibit inherently high computational parallelism, they are often compute-bound with larger FLOPs per token, which constrains their practical efficiency. Empirical results further shows that acceleration strategies such as dual cache and parallel decoding bring substantial speedups under small-batch settings. Overall, these insights are expected to encourage the community to deepen the exploration of DLM architectures and acceleration strategies, fostering a more diverse and efficient ecosystem.

### v3

@@ 5.2.2 Analysis of AR Models
In the decoding phase of AR models, the model generates only one token at each step, incurring per-token FLOPs of $\mathcal{O}(Bd^{2})+\mathcal{O}(BLd)$ and MOPs of $\mathcal{O}(d^{2})+\mathcal{O}(BLd)$ due to Key-Value (KV) cache reads. According to Figure 3, at the batch size of one, the arithmetic intensity of the AR decoding process is the lowest, making it a memory-bound workload. Consequently, its achievable throughput is mainly limited by memory bandwidth and can be asymptotically expressed as:
$\text{Throughput}\approx\begin{cases}\dfrac{\mathcal{O}(B_{\text{mem}}B)}{\mathcal{O}(d^{2})+\mathcal{O}(Ld)}=\dfrac{\mathcal{O}(B)}{\mathcal{O}(d^{2})+\mathcal{O}(Ld)},&L\ll d,\\[10.0pt] \dfrac{\mathcal{O}(B_{\text{mem}})}{\mathcal{O}(d^{2})+\mathcal{O}(Ld)}=\dfrac{\mathcal{O}(1)}{\mathcal{O}(d^{2})+\mathcal{O}(Ld)},&L\gg d.\end{cases}$ (5)
This formulation is applied in the analysis of generation length (with the batch size fixed at $1$) and batch size (under the condition $L\ll d$).
@@
Effect of Batch Size. In our evaluation setting (with generation length fixed at 256 tokens and relatively short prompts), the condition $L\ll d$ holds. In this regime, the denominator remains constant while the numerator scales linearly with batch size $B$ (because of the arithmetic intensity is $\mathcal{O}(B)$ ), leading to nearly linear throughput growth until the system reaches the compute roof. Once the arithmetic intensity reaches the compute roof, throughput remains constant as the workload becomes compute-bound.
@@ 5.2.3 Analysis of Diffusion Models
Therefore, the arithmetic intensity satisfies:
$\text{ArInt}\approx\begin{cases}\mathcal{O}(BL),&L\ll d,\\[6.0pt] \mathcal{O}(L),&L\gg d.\end{cases}$ (7)
Effect of Generation Length. For very short $L$ , the workload is still memory-bound. The throughput can be derived from Equation 7:
$\text{Throughput}\approx\frac{\mathcal{O}(B_{\text{mem}}BL)}{\mathcal{O}(BLd^{2})+\mathcal{O}(BL^{2}d)}=\frac{\mathcal{O}(1)}{\mathcal{O}(d^{2})+\mathcal{O}(Ld)}.$ (8)
Since $L\ll d$, the $\mathcal{O}(d^{2})$ term dominates, so the throughput is roughly stable with respect to $L$. However, in our experiments shown in Figure 3, the arithmetic intensity of the DLM quickly exceeds the ridge point, indicating that it soon operates in the compute-bound regime. In this case, the throughput is:
$\text{Throughput}\approx\frac{P_{\text{max}}}{\mathcal{O}(BLd^{2})+\mathcal{O}(BL^{2}d)}=\frac{P_{\text{max}}}{\mathcal{O}(Ld^{2})+\mathcal{O}(L^{2}d)},$ (9)
where $P_{\text{max}}$ represents the peak floating-point performance (Peak FLOPs). Here the numerator stays roughly constant, so as $L$ increases, the overall throughput decreases.
Effect of Batch Size. As shown in Figure 4, the DLM remains compute-bound (with $\text{ArInt}\approx\mathcal{O}(BL)$ ). Consequently, the throughput stays roughly flat as batch size $B$ increases, until it reaches the GPU memory limit. This trend is consistent with the observations in RQ1.
@@ 5.2.4 Analysis of Block Diffusion Models
In the decoding phase of block diffusion models, the sequence is partitioned into blocks of size G. Tokens within each block are generated in parallel via a diffusion process, incurring FLOPs of $\mathcal{O}(BGd^{2})+\mathcal{O}(BGLd)$. Despite the inter-block parallelism, the block-level progression follows an autoregressive scheme, benefiting from KV cache reuse. The resulting arithmetic intensity, as shown in Figure 3 and Figure 4, is inherently higher than that of standard AR models because the expensive KV cache reads are amortized across the parallel block generation. Consequently, its throughput can be formulated as:
$\text{Throughput}\approx\begin{cases}\dfrac{\mathcal{O}(B_{\text{mem}}BG)}{\mathcal{O}(Gd^{2})+\mathcal{O}(GLd)}=\dfrac{\mathcal{O}(BG)}{\mathcal{O}(Gd^{2})+\mathcal{O}(GLd)},&L\ll d,\\[10.0pt] \dfrac{\mathcal{O}(B_{\text{mem}}G)}{\mathcal{O}(Gd^{2})+\mathcal{O}(GLd)}=\dfrac{\mathcal{O}(G)}{\mathcal{O}(Gd^{2})+\mathcal{O}(GLd)},&L\gg d.\end{cases}$ (10)
Effect of Generation Length. Similar to AR, the block diffusion model is memory-bound when the batch size is 1. After simplification, its throughput is the same as that of the AR model, and thus follows the same trend: it remains roughly constant for short $L$, where the computation is dominated by $\mathcal{O}(d^{2})$ (Figure 1a-b) and decreasing as $L$ grows, where the complexity is dominated by $\mathcal{O}(Ld)$ (Figure 1c).
Effect of Batch Size. In our evaluation setting (with generation length fixed at 256 tokens and relatively short prompts), the condition $L\ll d$ holds. For block diffusion models, the arithmetic intensity is $\mathcal{O}(BG)$, which is higher than AR under the same conditions. As a result, block diffusion reaches the compute-bound regime at a smaller batch size than AR. As batch size increases, the throughput of block diffusion models behaves much like AR: throughput grows almost linearly with $B$ until it hits the compute roof (Figure 2a); beyond that point, throughput stays steady as the workload becomes compute-bound (Figure 2b).
@@ 5.3.1 Effect of sequence length
For block diffusion models, the “+parallel” strategy is highly effective when evaluated with a batch size of 1, consistently achieving the fastest decoding performance across all sequence lengths and even outperforming the AR baseline. We observe that this acceleration effect further amplifies as sequence length increases, reaching up to 3.1× speedup over the original configuration (i.e., 142.33 vs. 46.20 tokens/s). This improvement stems from the model’s ability to generate multiple high-confidence tokens simultaneously within each forward pass, effectively reducing the number of decoding steps and increasing overall throughput. Consequently, the effectiveness of different parallel strategies can be evaluated by their degree of parallel generation, which can be quantified by the number of tokens per forward step (TPF) [17-arxiv-2025-yuxin]. This metric directly reflects how efficiently a model converts computational intensity into output throughput, providing a unified lens for analyzing diffusion-based acceleration methods.
@@ 6 Future Directions
Inference Infrastructure for DLMs. Despite the impressive progress of DLMs, major machine learning ecosystems provide only limited optimization and deployment support for DLMs, posing practical challenges for researchers and developers. In particular, DLMs lack mature, open-source inference infrastructure akin to vLLM [18-sosp-2023-woosuk], making efficient serving of DLMs difficult.
@@
Applications of DLMs. Although DLMs are generally less efficient than AR models, their parallel architecture and iterative refinement can offer advantages in compute-rich or latency-sensitive scenarios. We highlight two representative application directions: (1) on-device inference. Combining with parallel decoding techniques, DLMs can achieve higher arithmetic intensity and throughput efficiency, making them well-suited for latency-sensitive and low-concurrency environments such as in-vehicle systems, mobile devices, and edge computing platforms. (2) structured and non-sequential generation. DLMs exhibit notable advantages in structured or non-sequential generation tasks such as code synthesis, table generation, and logical planning [19-arxiv-2025-samar, 20-arxiv-2025-chengze]. This suggests that the DLMs hold strong potential for scenarios requiring global semantic planning or multi-constraint generation.
@@ 7 Conclusion
In this report, we present a preliminary investigation into the efficiency of DLMs through both theoretical and empirical analyses. Our analysis reveals that existing evaluations remain limited in scope and consistency, lacking a comprehensive understanding of DLM efficiency across diverse conditions. In response, we introduce a theoretical analysis based on the model-side and hardware-size computational efficiency, revealing that while DLMs exhibit inherently high computational parallelism, they are often compute-bound with larger FLOPs per token, which constrains their practical efficiency. Empirical results further shows that acceleration strategies such as dual cache and parallel decoding bring substantial speedups under small-batch settings. Overall, these insights are expected to encourage the community to deepen the exploration of DLM architectures and acceleration strategies, fostering a more diverse and efficient ecosystem.
Appendix A Arithmetic Intensity Estimation
We estimate the specific expressions of arithmetic intensity for different model architectures, including AR, DLM, and block diffusion. The notations we will be using are as follows:
Notation Description
NN Total number of parameters.
$B$ Batch size.
$n_{l}$ Number of layers.
$n_{h}$ Number of attention heads.
$n_{d}$ Dimension of each attention head.
$\alpha$ Expansion ratio of the FFN layer.
$L$ Sequence length.
$d$ Hidden size.
$G$ Block size.
The mathematical estimation expressions for the arithmetic intensity of AR, Block Diffusion, and DLM are presented as follows:
$\text{ArInt}_{AR}=\frac{BN}{N+Bn_{l}n_{h}n_{d}L}.$ (11)
$\text{ArInt}_{BlockDiffusion}=\frac{2Bn_{l}(2Gd^{2}+\alpha^{2}Gd^{2}+LGd)}{N+2Bn_{l}dL+Bn_{l}dG}.$ (12)
$\text{ArInt}_{DLM}=\frac{2Bn_{l}(2Ld^{2}+\alpha^{2}Ld^{2}+L^{2}d)}{N+Bn_{l}dL}.$ (13)